# Quantifying the Dunkelflaute — An analysis of variable renewable energy droughts in Europe

### v3

@@ 4.5 Interconnection scenarios
Multi-regional VRE drought analyses based on VRE availability time series can be conducted based on two extreme assumptions of electricity transmission between countries [kittel_measuring_2024]: either perfect interconnection across all countries (“copper plate scenario”) or complete isolation of all countries (“island scenario”). For the former, all countries are treated as one single pan-European node. All regional time series are combined into a composite using capacity-weighted averages, with weights according to the capacity assumptions from TYNDP 2022. For the latter scenario, thresholds are scaled as discussed above. Accordingly, regional VRE portfolios are constructed using capacity-weighted averages of all contributing technologies. Note that more policy-relevant scenarios between these extreme cases, i.e., with limited interconnection capacity, cannot be applied to a purely time-series-based analysis without additional optimization or assumption-based heuristics.
@@
5 Data and code availability
The code and input data of the drought analysis tool are available in a public GitLab repository at https://gitlab.com/diw-evu/variable_renewable_energy_droughts_analyzer. The power sector model code and respective input data are provided in another public GitLab repository at https://gitlab.com/diw-evu/projects/quantifying_the_dunkelflaute_energy_system_analysis. High resolution, interactive, or animated versions versions of most figures presented in this article are available in a public Zenodo repository at https://doi.org/10.5281/zenodo.13843390.
@@ Acknowledgments
Quantifying the Dunkelflaute
An analysis of variable renewable energy droughts in Europe
@@
Cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau_{i}$ with $i\in[0.1,...,1]$ , sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. An interactive version of this Figure which allows zooming and rotating is available online [kittel_high_2025].
@@
Cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau_{i}$ with $i\in[0.1,...,1]$, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold $\tau_{0.75}$. An interactive version of this Figure which allows zooming and rotating is available online [kittel_high_2025].

### v4

@@ 4.5 Interconnection scenarios
Multi-regional VRE drought analyses based on VRE availability time series can be conducted based on two extreme assumptions of electricity transmission between countries [51]: either perfect interconnection across all countries (“copper plate scenario”) or complete isolation of all countries (“island scenario”). For the former, all countries are treated as one single pan-European node. All regional time series are combined into a composite using capacity-weighted averages, with weights according to the capacity assumptions from TYNDP 2022. For the latter scenario, thresholds are scaled as discussed above. Accordingly, regional VRE portfolios are constructed using capacity-weighted averages of all contributing technologies. Note that more policy-relevant scenarios between these extreme cases, i.e., with limited interconnection capacity, cannot be applied to a purely time-series-based analysis without additional optimization or assumption-based heuristics.
@@
5 Data availability
The input data of the drought analysis tool are available in a public GitLab repository at https://gitlab.com/diw-evu/variable_renewable_energy_droughts_analyzer. The power sector model input data are provided in another public GitLab repository at https://gitlab.com/diw-evu/projects/quantifying_the_dunkelflaute_energy_system_analysis. High resolution, interactive, or animated versions versions of most figures presented in this article are available in a public Zenodo repository at https://doi.org/10.5281/zenodo.13843390.
6 Code availability
The code of the drought analysis tool is available in a public GitLab repository at https://gitlab.com/diw-evu/variable_renewable_energy_droughts_analyzer. The power sector model code is provided in another public GitLab repository at https://gitlab.com/diw-evu/projects/quantifying_the_dunkelflaute_energy_system_analysis.
@@ Acknowledgments
Cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau_{i}$ with $i\in[0.1,...,1]$ , sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. An interactive version of this Figure which allows zooming and rotating is available online [52].
@@
Cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau_{i}$ with $i\in[0.1,...,1]$, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold $\tau_{0.75}$. An interactive version of this Figure which allows zooming and rotating is available online [52].